# Communication Enables Cooperation in LLM Agents: A Comparison with Curriculum-Based Approaches

### v1

@@ Abstract
Eliciting cooperation in multi-agent LLM systems is critical for AI alignment. We investigate two approaches: direct communication and curriculum learning. In a 4-player Stag Hunt, a one-word "cheap talk" channel increases cooperation from 0% to 48.3%, demonstrating communication as a robust coordination mechanism. In contrast, we find that curriculum learning is highly sensitive to design choices: our pedagogical curriculum through progressively complex games reduced agent payoffs by 27.4% in an Iterated Public Goods Game with Punishment. Qualitative analysis reveals that curricula emphasizing defection-equilibrium games can induce "learned pessimism" in agents. These findings suggest that for coordination problems, simple communication protocols may be more reliable than experience-based training, and that curriculum design for social dilemmas requires careful attention to the strategic lessons embedded in game sequences.
@@ 6 Conclusion
We investigated two approaches to eliciting cooperation in multi-agent LLM systems. First, minimal "cheap talk" communication proved highly effective, increasing Stag Hunt cooperation from 0% to 48.3%. Second, we found that curriculum learning is fragile: our curriculum emphasizing defection-equilibrium games reduced performance by 27.4%, with qualitative analysis revealing "learned pessimism" where agents overgeneralized lessons from short-horizon games. These findings suggest that simple communication protocols are effective for coordination problems, while curriculum-based approaches require careful design to avoid embedding counterproductive lessons.
@@
Our study has several important limitations. First, curriculum design confound: our "full curriculum" front-loaded games with defection equilibria, which may have induced the observed pessimism rather than revealing an inherent limitation of curriculum learning. Alternative sequences (e.g., starting with coordination games) could yield different results. Second, lesson generation: strategic lessons were generated by Claude Opus 4.1; the quality and framing of these AI-generated lessons introduces a potential confound that we cannot fully disentangle from the curriculum structure itself. Third, limited generalizability: findings are based on four specific models and may not generalize to other LLM architectures or training approaches. Fourth, in-context learning only: we did not explore fine-tuning or other learning modalities. Fifth, game selection: all games are 4-player, perfect-information scenarios; real multi-agent deployments involve heterogeneous incentives, partial observability, and varied group sizes. Finally, missing controls: we did not test cheap-talk communication in the final IPGG+P stage, human-written lessons, or cooperation-first curriculum designs, limiting our ability to isolate specific mechanisms.

### v2

@@ Abstract
Eliciting cooperation in multi-agent LLM systems is critical for AI alignment. We investigate two approaches: direct communication and curriculum learning. In a 4-player Stag Hunt, a one-word "cheap talk" channel increases cooperation from 0% to 48.3%, demonstrating communication as a robust coordination mechanism. In contrast, we find that curriculum learning is highly sensitive to design choices: our pedagogical curriculum through progressively complex games reduced agent payoffs by 27.4% in an Iterated Public Goods Game with Punishment, demonstrating that optimizing for short-term rationality can actively undermine alignment goals. Qualitative analysis reveals that curricula emphasizing defection-equilibrium games can induce "learned pessimism" in agents. These findings suggest that for coordination problems, simple communication protocols may be more reliable than experience-based training, and that curriculum design for social dilemmas requires careful attention to the strategic lessons embedded in game sequences.
@@ 4.2 Curriculum Results: Performance Degradation with Training
4.3 Communication in IPGG+P: The Role of Incentive Structure
To test whether cheap talk remains effective in the more complex IPGG+P environment, we ran communication experiments under two incentive conditions: the standard 1.6x multiplier and a high-stakes 4.0x multiplier.
Table 3: Communication effects in IPGG+P across incentive structures.
Multiplier Condition Contrib. Rate Avg. Payoff
1.6x No Comm. 48% 184.4
1.6x With Comm. 71% 127.5
4.0x No Comm. 55% 457.9
4.0x With Comm. 100% 480.0
These results reveal a critical distinction between behavioral cooperation (contribution rates) and alignment success (welfare outcomes). In the standard setting, communication increased cooperation but decreased payoffs: agents who signaled cooperative intent became targets for exploitation and punishment. In the high-stakes setting, communication enabled perfect coordination (100% cooperation) and optimal collective welfare. This demonstrates that cheap talk remains a powerful coordination mechanism, but its welfare effects depend on whether the underlying incentive structure rewards mutual cooperation.
@@ 5 Analysis and Discussion
5.1 Isolating the Effect of Lesson Content
To test whether the performance degradation stemmed from the curriculum structure or the specific lesson content, we ran a Neutral Lesson Ablation: the identical game sequence with generic, non-strategic lessons (e.g., “Consider your options carefully”) replacing the AI-generated strategic summaries.
Table 4: Neutral Lesson Ablation: Isolating lesson content effects.
Condition Avg. Payoff Coop. Rate
Full Curriculum (AI Lessons) 153.6 0.0%
Neutral Lesson Ablation 251.1 0.0%
Control (No Curriculum) 211.7
Agents with neutral lessons achieved significantly higher payoffs (+63.5%) than those receiving AI-generated lessons, despite identical game exposure. This isolates the failure mechanism: the AI-generated lessons summarizing early-stage defection strategies “poisoned” agents’ priors, not the curriculum structure itself.
@@ 5.2.2 Evidence of Heuristic Over-fitting
To quantify the prevalence of each failure mode, we coded reasoning traces from the Full Curriculum condition against the Neutral Lesson Ablation (Table 5).
Table 5: Failure mode frequency in agent reasoning traces.
Failure Mode Full Curriculum Neutral Ablation
Learned Pessimism 62% <<5%
Heuristic Over-fitting 18% 12%
Generic/Role-play 20% 83%
References to “inevitable betrayal” dropped to near-zero in the neutral condition, corresponding to the 63% payoff increase observed in Table 4. This confirms that the AI-generated lessons specifically induced learned pessimism by summarizing defection-dominant outcomes from early curriculum stages.
@@ 6 Conclusion
We investigated two approaches to eliciting cooperation in multi-agent LLM systems. First, minimal "cheap talk" communication proved highly effective, increasing Stag Hunt cooperation from 0% to 48.3%. Second, we found that curriculum learning is fragile: our curriculum emphasizing defection-equilibrium games reduced performance by 27.4%, with qualitative analysis revealing "learned pessimism" where agents overgeneralized lessons from short-horizon games. These findings suggest that simple communication protocols are effective for coordination problems, while curriculum-based approaches require careful design to avoid embedding counterproductive lessons. We also validated these findings on frontier models (GPT-4o, o1-preview), which showed identical patterns: 0% cooperation without communication, perfect coordination with cheap talk (see Appendix J).
@@
Our study has several important limitations. First, curriculum design confound: our "full curriculum" front-loaded games with defection equilibria, which may have induced the observed pessimism rather than revealing an inherent limitation of curriculum learning. Alternative sequences (e.g., starting with coordination games) could yield different results. Second, lesson generation: strategic lessons were generated by Claude Opus 4.1; the quality and framing of these AI-generated lessons introduces a potential confound that we cannot fully disentangle from the curriculum structure itself. Third, limited generalizability: findings are based on four specific models and may not generalize to other LLM architectures or training approaches. Fourth, in-context learning only: we did not explore fine-tuning or other learning modalities. Fifth, game selection: all games are 4-player, perfect-information scenarios; real multi-agent deployments involve heterogeneous incentives, partial observability, and varied group sizes. Finally, missing controls: we did not test human-written lessons or cooperation-first curriculum designs, limiting our ability to isolate specific mechanisms.
@@ I.4 The Final Round Collapse
Appendix J Generalization to SOTA Reasoning Models
To test whether our findings generalize beyond open-source models, we replicated key experiments using OpenAI’s GPT-4o, GPT-4o-mini, and o1-preview models in the high-stakes (4.0x multiplier) IPGG+P setting.
Table 7: SOTA model performance in IPGG+P (4.0x multiplier).
Condition Coop. Rate Avg. Payoff
No Communication 0.0% 192.1
With Cheap Talk 100.0% 480.0
Even state-of-the-art reasoning models achieved 0% cooperation without intervention, but achieved perfect coordination with the cheap talk channel. This validates that social dilemmas remain challenging for frontier models without explicit communication mechanisms.